my name is alexis and I am having troubles with my homework on problem solving such as 16- J= 12, what is the value of J? Our teacher taught us to make a positive 16 which crosses out the other 16 but = 28 when added to 12. Of course 4 is the answer, but how do I get to this answer on this specific problem. an example of what i'm doing is; s-38=57
38 38
------
0-95
s=95

Answer by elima(1433)

16- J= 12
What your teacher meant is doing the opposite operation to get the j by itself. So lets look at the left side: 16-j; this is a positive 16, so you will nedd to add it by the opposite of +16 which is -16, so you can make that a zero. And what you do to one side you do the exact same thing to the other side.So we have:
16-j+(-16)=12-16, the 16 cancels out on the left and you are left with:
-j=-4; but we do not want to leave the j as negative, so you will divide it by -1, and what you do to one side you do to the other. So you are left with:
j=4